# On the $\ell$-th largest degree of an intersecting family

### v1

@@ Abstract
Let $\mathcal{F}\subset\binom{[n]}{k}$ be an intersecting family. For an element $i\in[n]$, the degree of $i$ is the number of sets in $\mathcal{F}$ that contain $i$. Assume that the degrees are ordered as $d_{1}\geq d_{2}\geq\cdots\geq d_{n}$. Huang and Zhao showed that if $n>2k$, then the minimum degree satisfies $d_{n}\leq\binom{n-2}{k-2}$, with the maximum attained by the $1$-star. We strengthen this result by proving that for $n\geq 2k+1$, the $(2k+1)$-th largest degree satisfies $d_{2k+1}\leq\binom{n-2}{k-2}$, thereby confirming a conjecture of Frankl and Wang. Furthermore, we prove that for $k>50$ and $n>\frac{11}{2}k$, the $(k+2)$-th largest degree $d_{k+2}$ is already at most $\binom{n-2}{k-2}$. The techniques we developed also yield an tight upper bound for the $(\ell+1)$-th largest degree $d_{\ell+1}$ for $\ell\leq k$ and sufficiently large $n$.
@@ 1 Introduction
Let $k>50$, $n>\frac{11}{2}k$, and let $\mathcal{F}\subset\binom{[n]}{k}$ be an intersecting family. Then
@@
Next we consider the $\ell$-th largest degree of an intersection family, Frankl and Wang [4] have determined the maximum possible values for $d_{\ell}(\mathcal{F})$ for $2\leq\ell\leq 4$. For the $(\ell+1)$-degree problem when $\ell\geq 4$, inspired by the Hilton–Milner construction, they provided the following example.
@@
They also prove that for $4\leq\ell\leq k$, $d_{\ell+1}\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$ when $n>2\ell^{2}k$. In this paper, we obtain an improved range for large $\ell$ via the following theorem.
@@
Suppose $20\leq\ell\leq k$, and $n>\frac{k^{2}}{\ell}+64k$. Then for every intersecting family $\mathcal{F}\subset\binom{[n]}{k}$, we have
@@ 2 The shifting method
Theorem 2.5 ([1, 5]).
Suppose $\mathcal{F}_{1}\subset\binom{[n]}{a}$ and $\mathcal{F}_{2}\subset\binom{[n]}{b}$ are $t$-cross-intersecting families with $|\mathcal{F}_{1}|=x_{1}$ and $|\mathcal{F}_{2}|=x_{2}$. Let $\mathcal{L}(a,x_{1})$ denote the family consisting of the first $x_{1}$ $a$-sets in the lexicographic order, and similarly define $\mathcal{L}(b,x_{2})$. Then $\mathcal{L}(a,x_{1})$ and $\mathcal{L}(b,x_{2})$ are also $t$-cross-intersecting.
We note that the first $\binom{n-t}{a-t}$ $a$-sets in the lexicographic order on $\binom{[n]}{a}$ are exactly those containing the set $[t]$. Similarly, the first $\binom{n-t}{b-t}$ $b$-sets in the lexicographic order on $\binom{[n]}{b}$ are exactly those containing $[t]$. Moreover, when n>a+btn>a+b-t, for every set $A\in\binom{[n]}{a}$ that does not contain the whole set $[t]$, there exists a set $B$ with $t]\subset B\in\binom{[n]}{b}$ such that $|A\cap B|\leq t-1$. This observation yields the following corollary.
Corollary 2.6.
Under the same assumptions as in Theorem 2.5, if n>a+btn>a+b-t, then either $x_{1}\leq\binom{n-t}{a-t}$ or $x_{2}\leq\binom{n-t}{b-t}$.
@@ 4 Estimating the $(\ell+1)$-th largest degree
In this section, we first prove Theorem 1.4, which states that for $k$ sufficiently large and $n>5.5k$, any intersecting family $\mathcal{F}\subset\binom{[n]}{k}$ satisfies $d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$. As we mentioned earlier, one cannot hope for $d_{k+1}(\mathcal{F})\leq\binom{n-2}{k-2}$ due to the Hilton–Milner example. Therefore to have the $1$-star as the extremal example, the index $k+2$ is best possible. We then apply similar techniques to prove Theorem 1.6, which determines the maximum value of $d_{\ell+1}(\mathcal{F})$ when $k\geq\ell\geq 20$ and $n$ is sufficiently large.
For a fixed $\ell$ and a subset $S\subseteq[\ell]$, we define the degree of $S$ (with respect to $\mathcal{F}$) as the size of the family
@@
The following lemma shows that if two disjoint subsets $S_{1},S_{2}\subseteq[\ell]$ both have large degrees, a contradiction can be derived.
@@
Let $\mathcal{F}$ be an $\ell$-shifted $k$-uniform intersecting family, and let $S_{1},S_{2}\subseteq[\ell]$ be disjoint subsets. Then either
$d_{S_{1}}(\mathcal{F})\leq\binom{n-\ell}{k-\ell-1+|S_{2}|}$
@@
For $X\subseteq[\ell]$, define
$\mathcal{F}(X):=\{A\setminus X\mid A\in\mathcal{F},\ A\cap[\ell]=X\}.$
For $i=1,2$, the family $\mathcal{F}(S_{i})$ is a $(k-|S_{i}|)$-uniform family on the ground set $n]\setminus[\ell]$. Take arbitrary $A\in\mathcal{F}(S_{1})$ and $B\in\mathcal{F}(S_{2})$. By part (c) of Lemma 2.13, since $S_{1}$ and $S_{2}$ are disjoint, we have,
$|A\cap B|\geq\ell+1-|S_{1}|-|S_{2}|.$
Hence, $\mathcal{F}(S_{1})$ and $\mathcal{F}(S_{2})$ are $(\ell+1-|S_{1}|-|S_{2}|)$-cross-intersecting.
Set $t:=\ell+1-|S_{1}|-|S_{2}|$. Applying Corollary 2.6 with parameters $a=k-|S_{1}|$, $b=k-|S_{2}|$, $t=t$, and ground set size $n$-$\ell$, we obtain that either
$|\mathcal{F}(S_{1})|\leq\binom{(n-\ell)-t}{a-t}\quad\text{or}\quad|\mathcal{F}(S_{2})|\leq\binom{(n-\ell)-t}{b-t}$.
Since the binomial coefficient $\binom{x}{y}$ is increasing in $x$ for $x\geq y\geq 0$, and $(n-\ell)-t\leq n$-$\ell$, we can relax both upper bounds to
$|\mathcal{F}(S_{1})|\leq\binom{n-\ell}{a-t}=\binom{n-\ell}{k-\ell-1+|S_{2}|}$,
$|\mathcal{F}(S_{2})|\leq\binom{n-\ell}{b-t}=\binom{n-\ell}{k-\ell-1+|S_{1}|}$.
respectively. Since $d_{X}(\mathcal{F})=|\mathcal{F}(X)|$ for any $X\subseteq[\ell]$, the conclusion follows. ∎
@@
By Proposition 2.11, we may assume that $\mathcal{F}$ is $(k+2)$-shifted and that $d_{1}\geq d_{2}\geq\dots\geq d_{k+2}$ are the $k+2$ largest degrees. We are given the condition $n>\frac{11}{2}k$.
Set $m=\bigl\lfloor\frac{k}{2}\bigr\rfloor$. Define
$\mathcal{G}=\bigl\{A\subset[k+2]\mid|A|\leq m,\ d_{A}(\mathcal{F})>\tbinom{n-k-2}{m-3}\bigr\}$.
We claim that $\mathcal{G}$ is an intersecting family. Indeed, if there are two disjoint sets $A,B\in\mathcal{G}$, then Lemma 4.1 with $\ell=k+2$, $S_{1}=A$, $S_{2}=B$ yields either
$d_{A}(\mathcal{F})\leq\tbinom{n-k-2}{|B|-3}\quad\text{or}\quad d_{B}(\mathcal{F})\leq\tbinom{n-k-2}{|A|-3}$.
Since $|A|,|B|\leq m$ by definition of $\mathcal{G}$, we have $|A|-3,|B|-3\leq m-3$. Consequently, $\tbinom{n-k-2}{|B|-3}\leq\tbinom{n-k-2}{m-3}$ and $\tbinom{n-k-2}{|A|-3}\leq\tbinom{n-k-2}{m-3}$, because the binomial coefficient $\tbinom{N}{x}$ is increasing in $x$ for $0\leq x\leq N/2$ and m3m-3 is much smaller than (nk2)/2(n-k-2)/2 under the given condition $n>\frac{11}{2}k$. This contradicts the membership of $A$ and $B$ in $\mathcal{G}$. Hence $\mathcal{G}$ is intersecting.
Let $\overline{\mathcal{G}}:=\{A\subset[k+2]\mid A\supseteq B\text{ for some }B\in\mathcal{G}\}$ be the upward closure of $\mathcal{G}$. Then $\overline{\mathcal{G}}$ is also an intersecting family. Applying Lemma 3.2 with $n=k+2$ and $k$ replaced by $m$ (note that $2m<k+2$), we obtain an element $i\in[k+2]$ such that for every $u\leq m$,
$d_{i}^{(u)}(\overline{\mathcal{G}})\leq\binom{k}{u-2}$.
In particular, for $u=2$ we have $d_{i}^{(2)}(\overline{\mathcal{G}})\leq\binom{k}{0}=1$, therefore $d_{i}^{(2)}(\mathcal{G})\in\{0,1\}$. Below we discuss two cases according to the value of $d_{i}^{(2)}(\mathcal{G})$.
@@
We split the set in $\mathcal{F}$ that contains $i$ into the following three parts and estimate the size of each part separately.
$D_{1}=\{A|i\in A\in\mathcal{F},|A\cap[k+2]|>m\},$
$D_{2}=\{A|i\in A\in\mathcal{F},|A\cap[k+2]|\leq m,A\cap[k+2]\not\in\mathcal{G}\},$
$D_{3}=\{A|i\in A\in\mathcal{F},A\cap[k+2]\in\mathcal{G}\}.$
(a) For $D_{1}$, consider $A\in\mathcal{F}$ with $i\in A$ and $|A\cap[k+2]|\geq m+1$. For each $j$ with $m+1\leq j\leq k$, there are at most $\binom{k+1}{j-1}$ $j$-element subsets $X\subseteq[k+2]$ containing $i$. For each such $X$, we have $d_{X}(\mathcal{F})\leq\binom{n-k-2}{k-j}$. Hence the size of $|D_{1}|$ is bounded by
$|D_{1}|\leq\sum_{j=m+1}^{k}\binom{k+1}{j-1}\binom{n-k-2}{k-j}\leq\Bigl(\sum_{j=m+1}^{k}\binom{k+1}{j-1}\Bigr)\binom{n-k-2}{k-m-1}\leq 2^{k+1}\binom{n-k-2}{k-m-1}$.
(b) For $D_{2}$, consider $A\in\mathcal{F}$ with $i\in A$, $|A\cap[k+2]|\leq m$, and $A\cap[k+2]\notin\mathcal{G}$. There are at most $2^{k+1}$ sets $X\subseteq[k+2]$ containing $i$. By definition of $\mathcal{G}$, each such $X$ satisfies $d_{X}(\mathcal{F})\leq\binom{n-k-2}{m-3}$. Thus,
$|D_{2}|\leq 2^{k+1}\binom{n-k-2}{m-3}$.
(c) For $D_{3}$, consider $A\in\mathcal{F}$ with $i\in A$ and $A\cap[k+2]\in\mathcal{G}$. Since $d_{i}^{(j)}(\mathcal{G})\leq d_{i}^{(j)}(\overline{\mathcal{G}})\leq\binom{k}{j-2}$ for $j\leq m$, and for each $j$-element set $X\in\mathcal{G}$ we have $d_{X}(\mathcal{F})\leq\binom{n-k-2}{k-j}$, we obtain (using $d_{i}^{(2)}(\overline{\mathcal{G}})=0$)
$|D_{3}|\leq\sum_{j=3}^{m}\binom{k}{j-2}\binom{n-k-2}{k-j}\leq\sum_{j=3}^{k}\binom{k}{j-2}\binom{n-k-2}{k-j}=\binom{n-2}{k-2}-\binom{n-k-2}{k-2}$.
Now we show that, under the conditions $n\geq 5.5k$ and $k\geq 50$,
$\binom{n-k-2}{k-2}\geq 2^{k+2}\binom{n-k-2}{k-m-1}>|D_{1}|+|D_{2}|.$
First, observe that
$\frac{\binom{n-k-2}{k-2}}{\binom{n-k-2}{k-m-1}}=\frac{\prod_{j=n-2k+1}^{\,n-2k+m-1}j}{\prod_{j=k-m}^{\,k-2}j}=\prod_{j=0}^{m-2}\frac{n-2k+j+1}{k-m+j}.$
When $n>5.5k$, each factor satisfies
$\frac{n-2k+j+1}{k-m+j}\geq\frac{n-2k+m-1}{k-2}>4.$
For $k\geq 50,n\geq 5.5k$, the following product is minimized when $n=5.5k$ and $k=50$. Direct computation then gives
$\prod_{j=0}^{9}\frac{n-2k+j+1}{k-m+j}\geq\left.\prod_{j=0}^{9}\frac{n-2k+j+1}{k-m+j}\right|_{k=50,n=5.5k}\geq 60\cdot 4^{10}.$
Consequently,
$\frac{\binom{n-k-2}{k-2}}{\binom{n-k-2}{k-m-1}}$ $=\prod_{j=0}^{m-2}\frac{n-2k+j+1}{k-m+j}$
$\geq\Biggl(\prod_{j=0}^{9}\frac{n-2k+j+1}{k-m+j}\Biggr)\Biggl(\prod_{j=10}^{m-2}\frac{n-2k+j+1}{k-m+j}\Biggr)$
@@
$\geq 2^{k+2}.$
Finally, combining the three parts,
$d_{i}(\mathcal{F})=|D_{1}|+|D_{2}|+|D_{3}|<\binom{n-2}{k-2}$.
Since $i\in[k+2]$ and $d_{1},d_{2}\dots,d_{k+2}$ are the $k+2$ largest degrees, we have $d_{k+2}(\mathcal{F})\leq d_{i}(\mathcal{F})$. Hence,
$d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$.
@@
Using similar techniques, we give the proof of Theorem 1.6.
Proof of Theorem 1.6:
In the following, we define the degree of a subset $S\subset[\ell+1]$ as
$d_{S}(\mathcal{F})=\bigl|\{A\in\mathcal{F}\mid A\cap[\ell+1]=S\}\bigr|.$
Suppose $n>\dfrac{k^{2}}{\ell}+64k$. By Lemma 2.11, we may assume that $\mathcal{F}$ is $(\ell+1)$-shifted.
Set $m=\bigl\lfloor\frac{\ell}{2}\bigr\rfloor$ and define
$\mathcal{G}=\bigl\{A\subset[\ell+1]\mid|A|\leq m,d_{A}(\mathcal{F})>\tbinom{n-\ell-1}{k-\ell-2+m}\bigr\}$.
Then, by Lemma 4.1 (applied with $\ell$ replaced by $\ell+1$), the (not necessarily uniform) family $\overline{\mathcal{G}}$ is intersecting. Applying Lemma 3.2 with $n=\ell+1$ and $k$ replaced by $m$, we obtain an element $i\in[\ell+1]$ such that $d_{i}^{(j)}(\overline{\mathcal{G}})\leq\binom{\ell-1}{j-2}$ for all $j\leq m$.
@@
We consider the following two cases (if $d_{i}^{(2)}(\overline{\mathcal{G}})=1$, we assume without loss of generality that $\{1,i\}\in\overline{\mathcal{G}}$):
Case 1: $d_{i}^{(2)}(\mathcal{G})=0$ or $d_{\{1,i\}}(\mathcal{F})\leq\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$.

### v3

@@ Abstract
Let $\mathcal{F}\subset\binom{[n]}{k}$ be an intersecting family. For an element $i\in[n]$, the degree of $i$ is the number of sets in $\mathcal{F}$ that contain $i$. Assume that the degrees are ordered as $d_{1}\geq d_{2}\geq\cdots\geq d_{n}$. Huang and Zhao showed that if $n>2k$, then the minimum degree satisfies $d_{n}\leq\binom{n-2}{k-2}$, with the maximum attained by the $1$-star. We strengthen this result by proving that for $n\geq 2k+1$, the $(2k+1)$-th largest degree satisfies $d_{2k+1}\leq\binom{n-2}{k-2}$, thereby confirming a conjecture of Frankl and Wang. Furthermore, we prove that for large $k$ and $n>12k$, the $(k+2)$-th largest degree $d_{k+2}$ is already at most $\binom{n-2}{k-2}$. The techniques we developed also yield a tight upper bound for the $(\ell+1)$-th largest degree $d_{\ell+1}$ for $\varepsilon k\leq\ell\leq k$ and sufficiently large $n>C_{\varepsilon}k$.
@@ 1 Introduction
It is not hard to check that $d_{1}(\mathcal{F})=\binom{n-1}{k-1}-\binom{n-k-1}{k-1}$, and $d_{2}(\mathcal{F})=\cdots=d_{k+1}(\mathcal{F})=\binom{n-2}{k-2}+1$.
@@
There exist an absolute constant $C=12$ such that, for sufficiently large $k$ and for $n>Ck$, let $\mathcal{F}\subset\binom{[n]}{k}$ be an intersecting family. Then
@@
Next we consider the $\ell$-th largest degree of an intersecting family, Frankl and Wang [5] determined the maximum possible values for $d_{\ell}(\mathcal{F})$ for $2\leq\ell\leq 4$. For the $(\ell+1)$-degree problem when $\ell\geq 4$, inspired by the Hilton–Milner construction, they provided the following example.
@@
They also proved that for $4\leq\ell\leq k$, $d_{\ell+1}\leq\binom{n-2}{k-2}+\binom{n-\ell-1}{k-\ell}$ when $n>2\ell^{2}k$. In this paper, we obtain an improved range for large $\ell$ via the following theorem.
@@
For any constant $\varepsilon$, there is a constant $C_{\varepsilon}$ such that, for sufficiently large $k$ and for $\varepsilon k\leq\ell\leq k+1$, if $n>C_{\varepsilon}k$ and $\mathcal{F}\subset\binom{[n]}{k}$ is an intersecting family, then
@@ 4 Estimating the $(\ell+1)$-th largest degree
In this section, we first prove Theorem 1.4, which states that for $k$ sufficiently large and $n>12k$, any intersecting family $\mathcal{F}\subset\binom{[n]}{k}$ satisfies $d_{k+2}(\mathcal{F})\leq\binom{n-2}{k-2}$. As we mentioned earlier, one cannot hope for $d_{k+1}(\mathcal{F})\leq\binom{n-2}{k-2}$ due to the Hilton–Milner example. Therefore to have the $1$-star as the extremal example, the index $k+2$ is best possible. We then apply similar techniques to prove Theorem 1.6, which determines the maximum value of $d_{\ell+1}(\mathcal{F})$ when $\ell=\Theta(k)$ and $n$ is sufficiently large.
For a fixed $\ell$ and a subset $S\subset[\ell]$, we define the degree of $S$ (with respect to $\mathcal{F}$) as the size of the family
@@
We first prove that if $\mathcal{A}\subset\binom{[n]}{a}$ and $\mathcal{B}\subset\binom{[n]}{b}$ are $t$-cross-intersecting, then their sizes can be bounded as follows.
@@
Assume $n>3\max\{a,b\}$, $\mathcal{A}\subset\binom{[n]}{a},\mathcal{B}\subset\binom{[n]}{b}$ are non-empty $t$-cross-intersecting families. Then we have
$|\mathcal{A}|\leq 2^{b}\binom{n-b}{a-t},~~|\mathcal{B}|\leq 2^{a}\binom{n-a}{b-t}$.
Proof.
We only prove the inequality for $|\mathcal{A}|$, then $|\mathcal{B}|$ can be bounded using the same method. Fix any $B\in\mathcal{B}$. Every $A\in\mathcal{A}$ must satisfy $|A\cap B|\geq t$, hence
$|\mathcal{A}|$ $\leq\bigl|\{A\mid|A\cap B|\geq t,|A|=a\}\bigr|$
$=\sum_{i=t}^{a}\binom{b}{i}\binom{n-b}{a-i}\leq 2^{b}\binom{n-b}{a-t}$.
The last inequality holds since when $n>3a$ and $i\leq a$, $\binom{n}{i}$ is monotone in $i$. ∎
The following lemma shows that under the shifted assumption, if two disjoint subsets $S_{1},S_{2}\subseteq[\ell]$ both have large degree, then a contradiction can be derived.
Lemma 4.2.
Let $\mathcal{F}$ be an $\ell$-shifted $k$-uniform intersecting family on $[n]$ with $n\geq 4k+\ell$, and let $S_{1},S_{2}\subseteq[\ell]$ be disjoint subsets. Then either
$d_{S_{1}}(\mathcal{F})\leq 2^{k-|S_{2}|}\binom{n-\ell-k+|S_{2}|}{k-\ell+|S_{2}|-1}$
@@
For a set $S\subset[\ell]$, let
$\mathcal{F}(S):=\{A\backslash S|S\subset A\in\mathcal{F}\}.$
Then $d_{S}(\mathcal{F})=|\mathcal{F}(S)|$. By Lemma 2.11 c), for each set $A\in\mathcal{F}(S_{1})$, $B\in\mathcal{F}(S_{2})$, we have $|A\cap B|\geq\ell+1-|S_{1}|-|S_{2}|$, therefore $\mathcal{F}(S_{1})\subset\binom{[n-\ell]}{k-|S_{1}|}$ and $\mathcal{F}(S_{2})\subset\binom{[n-\ell]}{k-|S_{2}|}$ are $(\ell+1-|S_{1}|-|S_{2}|)$-cross-intersecting.
If $\mathcal{F}(S_{2})$ is empty, then $d_{S_{2}}(\mathcal{F})=0$ so the statement holds. If $\mathcal{F}(S_{2})$ is non-empty, then by Lemma 4.1,
$|\mathcal{F}(S_{1})|$ $\leq 2^{k-|S_{2}|}\binom{n-\ell-k+|S_{2}|}{k-|S_{1}|-(\ell+1-|S_{1}|-|S_{2}|)}$
$=2^{k-|S_{2}|}\binom{n-\ell-k+|S_{2}|}{k-\ell-1+|S_{2}|}$.
Hence the statement is true. ∎
Corollary 4.3.
Let $\mathcal{F}$ be an $\ell$-shifted $k$-uniform intersecting family. Suppose $m=\lfloor\frac{\ell-1}{2}\rfloor$, $n\geq 4k+\ell$, and define $\mathcal{G}$ as the following family of subsets of $\ell$]:
$\mathcal{G}:=\bigl\{S\subset[\ell]\mid|S|\leq m,d_{S}(\mathcal{F})\geq 2^{k-m}\tbinom{n-k-\ell+m}{k-\ell+m}\bigr\}$.
Then $\mathcal{G}$ is intersecting.
Proof.
Let $f(j)=2^{k-j}\binom{n-k-\ell+j}{k-\ell+j-1}$. Note that when $n>3k$, for $j\leq m\leq\ell$, we have
$f(j+1)$ $=f(j)\cdot\frac{\binom{n-k-\ell+j+1}{k-\ell+j}}{2\binom{n-k-\ell+j}{k-\ell+j-1}}=f(j)\cdot\frac{n-k-\ell+j+1}{2(k-\ell+j)}\geq f(j).$
If $S_{1},S_{2}\in\mathcal{G}$ are disjoint, then by Lemma 4.2, we have either
$d_{S_{1}}(\mathcal{F})\leq 2^{k-|S_{2}|}\binom{n-\ell-k+|S_{2}|}{k-\ell+|S_{2}|-1}$
or
$d_{S_{2}}(\mathcal{F})\leq 2^{k-|S_{1}|}\binom{n-\ell-k+|S_{1}|}{k-\ell+|S_{1}|-1}$.
Without loss of generality, we assume that $d_{S_{1}}(\mathcal{F})\leq 2^{k-|S_{2}|}\binom{n-k-\ell+|S_{2}|}{k-\ell-1+|S_{2}|}$, then by the monotonicity of $f(j)$, we have
$d_{S_{1}}(\mathcal{F})$ $\leq 2^{k-|S_{2}|}\binom{n-k-\ell+|S_{2}|}{k-\ell-1+|S_{2}|}=f(|S_{2}|)\leq f(m)$
$=2^{k-m}\binom{n-k-\ell+m}{k-\ell+m-1}<2^{k-m}\binom{n-k-\ell+m}{k-\ell+m}$.
This contradicts the assumption $S_{1}\in\mathcal{G}$. ∎
@@
Let $m=\lfloor(k+1)/2\rfloor$. By Proposition 2.9, we may assume that $\mathcal{F}$ is $(k+2)$-shifted and that $d_{1}\geq d_{2}\geq\dots\geq d_{k+2}$ are the $k+2$ largest degrees. For a subset $X\subset[k+2]$, we define the degree of $X$ in $\mathcal{F}$ as
$d_{X}(\mathcal{F})=\bigl|\{A\in\mathcal{F}\mid A\cap[k+2]=X\}\bigr|.$
We are given the condition $n>12k$. Let $\mathcal{G}$ be the following family:
$\mathcal{G}:=\bigl\{S\subset[k+2]\mid|S|\leq m,d_{S}(\mathcal{F})\geq 2^{k-m}\tbinom{n-2k-2+m}{m-2}\bigr\}$.
By Corollary 4.3 with $\ell=k+2$, $\mathcal{G}$ is an intersecting family in $2^{[k+2]}$. Its upward closure $\overline{\mathcal{G}}$ is also intersecting. Therefore Lemma 3.2 shows there exists some element $i\in[k+2]$ such that $d^{(j)}_{i}(\overline{\mathcal{G}})\leq\binom{k}{j-2}$ for $j=1,2,\dots,m$, in particular, $d^{(2)}_{i}(\overline{\mathcal{G}})\in\{0,1\}$. Below we discuss two cases according to the value of $d_{i}^{(2)}(\overline{\mathcal{G}})$.
@@
We split the sets in $\mathcal{F}$ that contain $i$ into the following two parts and estimate the size of each part separately:
$D_{1}=\big\{A\big|i\in A\in\mathcal{F},A\cap[k+2]\in\mathcal{G}\big\},$
$D_{2}=\big\{A\big|i\in A\in\mathcal{F},A\cap[k+2]\not\in\mathcal{G}\big\}.$
We first estimate $|D_{1}|$:
$|D_{1}|$ $=\sum_{i\in X\in\mathcal{G}}d_{X}(\mathcal{F})\leq\sum_{j\geq 2}d_{i}^{(j)}(\mathcal{G})\binom{n-k-2}{k-j}\leq\sum_{j\geq 2}d_{i}^{(j)}(\overline{\mathcal{G}})\binom{n-k-2}{k-j}$
$\leq\sum_{j\geq 3}\binom{k}{j-2}\binom{n-k-2}{k-j}=\binom{n}{k-2}-\binom{n-k-2}{k-2}$.
The first inequality is because for fixed $X\in\mathcal{G}$, the number of $k$-set AA with $A\cap[k+2]=X$ is at most $\binom{n-k-2}{k-|X|}$. The second inequality follows from $\mathcal{G}\subset\overline{\mathcal{G}}$. The third inequality follows from Lemma 3.2 and $d^{(2)}_{i}(\overline{\mathcal{G}})=0$.
Next we estimate $|D_{2}|$. We prove the following claim: for $X\notin\mathcal{G}$, we have $d_{X}(\mathcal{F})\leq 2^{k-m}\binom{n-k-2}{k-m}$.
First suppose $|X|\leq m$, since $2m\leq k+1<k+2$, i.e. km>m2k-m>m-2, by the definition of $\mathcal{G}$, we have
$d_{X}(\mathcal{F})\leq 2^{k-m}\binom{n-2k-2+m}{m-2}\leq 2^{k-m}\binom{n-k-2}{k-m}$.
If $|X|>m$, then
$d_{X}(\mathcal{F})\leq\binom{n-k-2}{k-m}\leq 2^{k-m}\binom{n-k-2}{k-m}$.
Using the claim, $|D_{2}|$ can be bounded by
$|D_{2}|$ $=\sum_{X\not\in\mathcal{G}}d_{X}(\mathcal{F})\leq\sum_{X\subset[k+2]}2^{k-m}\binom{n-k-2}{k-m}\leq 2^{2k+2-m}\binom{n-k-2}{k-m}$.
Next we show that for sufficiently large $k$, when $n>12k$, we have
$2^{2k+2-m}\binom{n-k-2}{k-m}\leq\binom{n-k-2}{k-2}$.
Recall that $k/2\leq m\leq(k+1)/2$, thus
@@
When $k$ is sufficiently large,
[…]
This implies that
$d_{i}=|D_{1}|+|D_{2}|\leq\binom{n-2}{k-2}-\binom{n-k-2}{k-2}+\binom{n-k-2}{k-2}\leq\binom{n-2}{k-2}$.
which completes the proof of Case 1.
@@
Using similar techniques, we proof Theorem 1.6.
Proof of Theorem 1.6.
Let $m=\lfloor\frac{\ell}{2}\rfloor$. By Proposition 2.9, we can assume that $\mathcal{F}$ is $(\ell+1)$-shifted. Throughout the proof, we define the degree $d_{X}$ of a subset $X\subset[\ell+1]$(with respect to $\mathcal{F}$) by
$d_{X}(\mathcal{F}):=\bigl|\{A\in\mathcal{F}\mid A\cap[\ell+1]=X\}\bigr|.$
Denote by $\mathcal{G}$ the following set family:
$\mathcal{G}:=\big\{X\subset[\ell+1]\big||X|\leq m,d_{X}(\mathcal{F})\geq 2^{k}\binom{n-k}{k-m}\big\}$
Since $2^{k}>2^{k-m},k-m\geq k-\ell-1+m$, and $n-k\geq n-k-\ell-1+m$, when $n>5k$ the monotonicity of binomial coefficients yields
$2^{k}\binom{n-k}{k-m}\geq 2^{k-m}\binom{n-k-\ell-1+m}{k-\ell-1+m}$,
@@
so by Corollary 4.3, $\mathcal{G}$ is an intersecting family on $2^{[\ell+1]}$, then by Lemma 3.2, there is an $i\in[\ell+1]$ such that $d_{i}^{(j)}(\overline{\mathcal{G}})\leq\binom{\ell-1}{j-2}$ for $1\leq j\leq m$. If $d_{i}^{(2)}(\overline{\mathcal{G}})=1$, without loss of generality, we may assume that $\{1,i\}\in\overline{\mathcal{G}}$. Next we consider two cases according to the value of $d_{\{1,i\}}(\overline{\mathcal{G}})$.
Case 1: $d_{i}^{(2)}(\overline{\mathcal{G}})=0$ or $d_{\{1,i\}}(\mathcal{F})\leq\binom{n-\ell-1}{k-2}-\binom{n-k-2}{k-2}$.